one's intelligence quotient, or IQ, varies directly as a person's mental age and inversely as that person's chronological age. A person with a mental age of 27 and a chronological age of 20 has an IQ of 135. What is the chronological age of a person with a mental age of 30 and an IQ of 120?
Answer by sudhanshu_kmr(1152)


IQ * CHRONOLOGIAL AGE / MENTAL AGE = CONSTANT
135*20/27 = 120*X/30
=> X = 135 *20*30 /(120*27)
=> X = 25

so, chronologial age = 25


Everyone is welcome to get free help online: sudhanshu.cochin@gmail.com